OpenStudy (anonymous):

0.00000003937 in scientific notation.

OpenStudy (anonymous):

3.937x10^-8

OpenStudy (doc.brown):

\[0.00000003937 =\]\[0.00000003937\times 1\]\[0.0000003937\times 10^{-1}\]\[0.000003937\times 10^{-2}\]\[0.00003937\times 10^{-3}\]\[0.0003937\times 10^{-4}\]\[0.003937\times 10^{-5}\]\[0.03937\times 10^{-6}\]\[0.3937\times 10^{-7}\]\[3.937\times 10^{-8}\]

OpenStudy (anonymous):

3.937x10^8